OpenStudy (mwhit2101):

It costs $50 plus $10 per hour to rent a park pavilion. Find the cost of renting the pavilion for 8 hours.

OpenStudy (mwhit2101):

130 150 90 140

OpenStudy (awy):

so $50 is your base cost + $10 per hour

OpenStudy (awy):

so, $50+$10(8 hours) = $130

OpenStudy (mwhit2101):

What?

OpenStudy (mwhit2101):

50+10*8=240..I dont see how you got 130

OpenStudy (awy):

oops, my bad, i forgot to add the 80 on my calculator

OpenStudy (anonymous):

50+80=130 so $50+10 *8

OpenStudy (awy):

wait no, its 130, do your order of operation correctly

OpenStudy (mwhit2101):

What...

OpenStudy (mwhit2101):

Very Confused....

OpenStudy (anonymous):

wat? i just said 130

OpenStudy (anonymous):

and the other stuff was showing my work

OpenStudy (mwhit2101):

How did you get that??????

OpenStudy (anonymous):

how i got the answer

OpenStudy (mwhit2101):

Why wouldn't it be 240 because like 50+10=60 and 60*8 is 240!!! THat makes the most sense

OpenStudy (awy):

becuase you didnt do your order of operation correctly

OpenStudy (awy):

multiply the 8 and 10 first then add the 50 dollars

OpenStudy (anonymous):

this is how i got it at the bottom